Solve -2|h - 7| = - 28
A. No solutions
B. h+21
C. h=-7, h=21
D. h=7, h= -21

Answer by stanbon(75887)

-2|h - 7| = - 28
-----------------------------
|h-7| = 14
-------
h-7 = 14 or h-7 = -14
h = 21 or h = -7
========================
Cheers,
Stan H.
========================

A. No solutions
B. h+21
C. h=-7, h=21
D. h=7, h= -21